OpenStudy (anonymous):

Use completing the square to solve the equation x^2 - 10x = -8. Which equation is used in the process? a) x2- 10x + 25 = 17 b) x2- 10x + 25 = -8 c) x2- 10x - 25 = -33

OpenStudy (anonymous):

x2- 10x + 25 = 17

OpenStudy (anonymous):

howd you get that

OpenStudy (anonymous):

I was going with the first one as well..

OpenStudy (anonymous):

u had to add (b/2)^2 in this case (10/2)^2

OpenStudy (anonymous):

Given that \[x^2 - 10x = -8. \rightarrow x^2 - 2 \times5 \times x = -8.\] comparing the middele term with 2ab we find a=x and b=5 since square of x is already given sq we need to add square of 5 on both sides so \[x^2 - 2 \times5 \times x + 5^2= 5^2-8.\] \[x^2 - 10x + 25= 25-8.\rightarrow x^2 - 10x + 25= 17\]

OpenStudy (anonymous):

a) x2- 10x + 25 = 17 ==> x2- 10x + 25- 17 ==> x2- 10x +8 ==> x2- 10x = -8

OpenStudy (anonymous):

okay! thank you!

OpenStudy (anonymous):

simplifying